Hi, I really need some help!
Write a quadratic equation with -7 and 5 as its roots.
Thank You!!!

Answer by stanbon(75887)

If -7 is a root then (x+7) is a factor.
If 5 is a root then (x-5) is a factor.
Therefore f(x)=(x-5)(x+7)=x^2+2x-35
is a quadratic funciton with roots -7 and 5.
Cheers,
stan H.